OpenStudy (anonymous):

A small planet has a spherical shape, its radius is 20 km, it is covered with ice whose thickness is 1/4 km (250 m). use differentials to approximate the volume of the ice. Please show a detailed solution, @Calculus1

OpenStudy (anonymous):

Dy=F'( x)(Dx) \[F'(x)=4\pi r^2\] Dx=1/4 \[Dy=4*\pi*20^2*.25--->1256km\] That's the volume of the Ice

OpenStudy (anonymous):

You have to take the derivative of the volume